# Empowering Multi-Robot Cooperation via Sequential World Models

### v1

@@ Abstract
Model-based reinforcement learning (MBRL) has shown significant potential in robotics due to its high sample efficiency and planning capability. However, extending MBRL to multi-robot cooperation remains challenging due to the complexity of joint dynamics. To address this, we propose the Sequential World Model (SeqWM), a novel framework that integrates the sequential paradigm into model-based multi-agent reinforcement learning. SeqWM employs independent, sequentially structured agent-wise world models to decompose complex joint dynamics. Latent rollouts and decision-making are performed through sequential communication, where each agent generates its future trajectory and plans its actions based on the predictions of its predecessors. This design enables explicit intention sharing, enhancing cooperative performance, and reduces communication overhead to linear complexity. Results in challenging simulated environments (Bi-DexHands and Multi-Quad) show that SeqWM outperforms existing state-of-the-art model-free and model-based baselines in both overall performance and sample efficiency, while exhibiting advanced cooperative behaviors such as predictive adaptation and role division. Furthermore, SeqWM has been successfully deployed on physical quadruped robots, demonstrating its effectiveness in real-world multi-robot systems. Demos and code are available at: https://github.com/zhaozijie2022/seqwm-marl.
@@ 1 Introduction
Model-based reinforcement learning (MBRL) has been widely applied to robotic systems due to its high sample efficiency (Wu et al., 2023) and planning capability (Sun et al., 2023). However, extending MBRL to multi-robot cooperation remains challenging. Early decentralized model-based multi-agent reinforcement learning (MARL) approaches built independent world models for each agent (Egorov & Shpilman, 2022), overlooking inter-agent couplings and hindering coordination. More recent centralized methods, by contrast, assume full observability or unrestricted communication (Chai et al., 2024; Toledo, 2024; Liu et al., 2024b), performing dynamics modeling and policy optimization in the joint space. These methods face challenges related to modeling complexity and communication overhead in robotic systems with high-dimensional observation and action spaces, limiting their deployment in real-world scenarios.
Between centralized and decentralized paradigms, the distributed sequential paradigm has rapidly developed in recent years and demonstrated unique advantages (Khan, 2025). It reformulates multi-agent decision-making as an autoregressive process: agents communicate and act in a certain order, with each updating its policy conditioned on messages and actions from predecessors (Wen et al., 2022; Hu et al., 2025). This design enables more consistent joint reasoning (Ding et al., 2024) and finer-grained credit assignment (Kuba et al., 2022; Wang et al., 2023) without relying on full communication. It also reduces communication overhead from $O(n^{2})$ in centralized settings to $O(n)$ in chain structures, while offering greater robustness under packet loss or disturbances.
Motivated by these advantages, we propose the Sequential World Model (SeqWM), which integrates the sequential paradigm into model-based MARL to structurally decompose dynamics modeling and action planning. For trajectory prediction, the joint dynamics are modeled as sequential agent-wise rollouts along the communication order, where each agent maintains an independent world model and conditions on the predicted trajectories and actions of its predecessors. For action planning, each agent performs multi-step lookahead based on its predecessors’ predictions, thereby preserving cooperative performance while restricting the search to a low-dimensional subspace consistent with the sequential structure.
Figure 1: Behaviors Comparison.
Methods Overhead
MAT (Hu et al., 2025) ndnd
MAZero (Liu et al., 2024b) $n^{2}Hd$
MARIE (Zhang et al., 2025a) $n^{2}KHd$
SeqComm (Ding et al., 2024) $2n^{2}Kd$
SeqWM nHdnHd
Table 1: Communication overhead.222  $n$ is agents number, $d$ is the message dimension, $H$ is prediction horizon, and $K$ is planning iterations.
We evaluate SeqWM in two challenging multi-robot cooperation environments: Bi-DexHands (Chen et al., 2024a) and Multi-Quad (Xiong et al., 2024), and further validate it on physical multi-robot cooperation tasks using two Unitree Go2-W robots. As shown in Figure 1 and Table 2, SeqWM achieves superior cooperative behaviors under comparable or even lower communication overheads. The key highlights of SeqWM are as follows:
@@
    By integrating the sequential paradigm, SeqWM structurally decomposes joint dynamics and planning, reducing modeling complexity and communication overhead;
@@
    Through explicit intention sharing, SeqWM enables the emergence of more advanced cooperative behavior patterns;
@@
    SeqWM outperforms current state-of-the-art methods on standard benchmarks and has been successfully deployed on physical robots.
These results collectively demonstrate that SeqWM, through sequential world modeling and planning, provides an effective pathway for multi-robot cooperation, balancing performance, efficiency, and real-world applicability.
@@ 2 Related Work
Model-based RL. In single-agent robotics, model-based RL has shown remarkable success due to its high sample efficiency (Zhou et al., 2024; Lancaster et al., 2024), with approaches such as PlanCP (Sun et al., 2023) and GPC (Qi et al., 2025) leveraging learned dynamics models to predict trajectories and optimize actions for physical robots. In contrast, existing model-based MARL methods often rely on centralized paradigms, hindering their practical deployment in multi-robot systems. For example, MACD (Chai et al., 2024) and CoDreamer (Toledo, 2024) use transformers or GNNs to integrate full state-action across all agents. Recent efforts such as MARIE (Zhang et al., 2025a) explored decentralized dynamics modeling, but still require communication at each prediction step for agent-wise aggregation. Different from these works, SeqWM assigns each agent an independent world model and predicts trajectories sequentially, significantly reducing modeling complexity and communication overhead for real-world deployment.
Sequential Paradigm. Recent studies in MARL have highlighted the advantages of the sequential paradigm (Khan, 2025), which enables fine-grained credit assignment (Kuba et al., 2022), efficient dynamics modeling (Zhang et al., 2025b), and scalable coordination (Xu et al., 2025). For example, MAT (Wen et al., 2022) and PMAT (Hu et al., 2025) model the decision-making process as a sequence prediction problem, employing transformers to autoregressively predict each agent’s actions. A2PO (Wang et al., 2023) and HARL (Zhong et al., 2024; Liu et al., 2024a) further introduce the sequential update scheme that bring clearer interpretability and ensure monotonic improvement. SeqComm (Ding et al., 2024) extends this idea to the communication domain, where agents exchange information in a sequential order, effectively mitigating non-stationarity while reducing communication overhead. Motivated by these benefits, we integrate the sequential paradigm into model-based MARL, leveraging world models to enhance planning and coordination.
@@ 3 Preliminaries
Problem Formulation. We model the fully cooperative task as a decentralized partially observable Markov decision process (dec-POMDP), $\mathcal{M}=\langle\mathcal{I},\mathcal{S},\mathcal{O},\mathcal{A},\Omega,\mathcal{P},\mathcal{R},\gamma\rangle$, where $\mathcal{I}=\{v^{1},\dots,v^{n}\}$ is the set of agents, $\mathcal{S}$ is the global state space, $\mathcal{O}=\prod_{i=1}^{n}O^{i}$ is the joint observation space, and $\mathcal{A}=\prod_{i=1}^{n}A^{i}$ is the joint action space. The observation function $\Omega:\mathcal{S}\times\mathcal{I}\rightarrow\mathcal{O}$ defines each agent’s perception of the environment, while the transition function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ specifies the environment dynamics. The reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ provides a shared scalar signal, and $\gamma$ is the discount factor. Each agent $v^{i}$ learns a local policy $\pi^{i}:O^{i}\rightarrow A^{i}$, which maps its observation $o^{i}$ to an action $a^{i}$. The objective is to learn a joint policy $\pi=\prod_{i=1}^{n}\pi^{i}$ that maximizes the expected discounted return $\sum_{\tau=t}^{\infty}\gamma^{\tau}r_{\tau}$.
@@ 4 Methodology
In this section, we propose SeqWM, a Sequential World Model for multi-robot cooperation. Unlike existing centralized or decentralized multi-agent world models, SeqWM decomposes the joint dynamics into agent-wise models arranged in a sequence. This design substantially reduces modeling complexity and communication overhead, enabling deployment in real-world scenarios.
@@ 4.1 Sequential Multi-Agent World Model
Multi-Agent World Model. As noted in Introduction, multi-robot cooperative tasks involve high-dimensional observation and action spaces, making it unsuitable to use reconstructing raw observations as the learning objective of the world model. Following (Hafner et al., 2025), we remove the explicit decoder and instead perform dynamics prediction entirely in a latent space. To facilitate distributed deployment, each agent maintains an independent world model without parameter sharing. Let $z^{i}_{t}$ denote the latent state of agent $v^{i}$ at timestep $t$, the world model can be defined as follows:
@@
All modules in SeqWM are implemented using MLPs, ensuring architectural simplicity and consistency. For communication function, we adopt concatenation operator $\oplus$ to facilitate modular training.333 We implement the concatenation using a mask-based scheme to ensure consistent dimensionality, with details provided in Appendix A.1. As shown in Section 5.4, this concise design achieves more stable training performance compared to alternatives such as cross-attention and recurrent neural networks (RNNs).
@@
where $\theta=\{\theta_{E},\theta_{D},\theta_{R},\theta_{Q}\}$, $H$ is the prediction horizon, $\lambda\in(0,1]$ is a constant that balances the contribution of each rollout step, $r_{t}$ is the ground-truth reward, $G_{t}$ is the TD target, and $\hat{z}^{i}_{t+1}=D^{i}\left(z^{i}_{t},a^{i}_{t},e^{i}_{t}\right)$ is the predicted latent state. The loss can be backpropagated to the encoder via $z^{i}_{t}=E^{i}\left(o^{i}_{t}\right)$, so do the dynamics and reward. The latent target $z^{i}_{t+1}=E^{i}\left(o^{i}_{t+1}\right)$ is detached with the stop-gradient operator $\textrm{sg}(\cdot)$ to prevent cyclic gradient flow and gradient degradation. Soft Cross-Entropy loss is used to match the discretized reward and Q-value predictions, with details provided in Appendix B.3. Additionally, to ensure modularity and scalability, each agent’s world model is trained independently, and the loss can not be backpropagated through the communication channel.
Based on Eq. (3), the encoder learns a compact latent space, while the dynamics and reward predictors minimize prediction errors in this space, ensuring alignment with real environment dynamics. The actor generates initial action estimates in the latent space to warm-start planning and is trained using the Heterogeneous-Agent Soft Actor-Critic (HASAC) (Haarnoja et al., 2018; Liu et al., 2024a) algorithm:
@@
where $\hat{a}^{i}_{t}=\pi^{i,\textrm{Act}}\left(z^{i}_{t},e^{i}_{t}\right)$, $\alpha$ is the entropy coefficient, and $\mathcal{H}[\cdot]$ denotes the entropy function.
@@ 4.2 Planning with Sequential Communication
Although Eq. (2) includes an actor, it does not serve directly as an explicit decision policy; instead, it provides initial action estimates for the planner. We next propose a sequential multi-agent planner based on Model Predictive Path Integral (MPPI) (Williams et al., 2015) that leverages the predictions of world models to optimize each agent’s action. In this framework, the actor contributes by narrowing the action search space to promising regions, while the planner ensures robust long-term decision-making and corrects suboptimal proposals from the actor. 444 Appendix C.3 provides an ablation analysis illustrating the complementary roles of the actor and planner.
@@
$V^{i}_{t+H}=\gamma^{H}Q^{i}(\hat{z}^{i}_{t+H},a^{i}_{t+H},e^{i}_{t+H})$ (5)
$+\sum_{h=t}^{t+H-1}\gamma^{h-t}R^{i}(\hat{z}^{i}_{h},a^{i}_{h},e^{i}_{h}),$
where $V^{i}_{t+H}$ represents the value estimate of a sampled action sequence, computed as the sum of predicted rewards over the horizon plus the terminal value given by the critic. Then, candidate sequences are ranked according to their evaluated values, and the highest-scoring subset is selected as the elite set. The action distribution is updated toward the statistics of these elite trajectories, thereby concentrating future sampling around high-value regions and progressively refining the search space. 555 The details of the planner and its implementationare provided in Appendix A.1 and  B.
@@
Low-pass Action Smoothing. To prevent mechanical wear caused by abrupt action changes during real-world execution, we integrate a low-pass filtering strategy (Kicki, 2025) into the planner. In each planning iteration, sampled action sequences are filtered along the temporal dimension to suppress high-frequency fluctuations. This smoothing enforces gradual action transitions across timesteps, reducing control discontinuities and promoting stable, consistent behavior on physical robots. Further details and formulations are provided in Appendix A.2.
Heuristic Early-Stopping. Considering the computational constraints of real robotic platforms, we design a motion-planning heuristic that terminates iterations when the Kullback-Leibler (KL) (Kullback & Leibler, 1951) divergence between consecutive action distributions falls below a threshold. This early-stopping criterion mitigates diminishing returns (Kobilarov, 2012), reducing computation while preserving plan quality. Further details and experiments regarding this design are provided in Appendix C.4.

### v2

@@ Abstract
Model-based reinforcement learning (MBRL) has shown significant potential in robotics due to its high sample efficiency and planning capability. However, extending MBRL to physical multi-robot cooperation remains challenging due to the complexity of joint dynamics and the reliance on synchronous communication. To address this, we propose the Sequential World Model (SeqWM), a novel framework that integrates the sequential paradigm into model-based multi-agent RL. SeqWM employs independent, autoregressive agent-wise world models to represent joint dynamics, where each agent generates its future trajectory and plans its actions based on the predictions of its predecessors. This design lowers modeling complexity, alleviates the reliance on communication synchronization, and enables the emergence of advanced cooperative behaviors through explicit intention sharing. Experiments in challenging simulated environments (Bi-DexHands and Multi-Quad) demonstrate that SeqWM outperforms existing state-of-the-art model-based and model-free baselines in both overall performance and sample efficiency, while exhibiting advanced cooperative behaviors such as predictive adaptation, temporal alignment, and role division. Furthermore, SeqWM has been successfully deployed on physical quadruped robots, validating its effectiveness in real-world multi-robot systems. Demos and code are available at: SeqWM-MARL.
@@ 1 Introduction
Model-based reinforcement learning (MBRL) has been widely applied to robotic systems due to its high sample efficiency (Wu et al., 2023) and planning capability (Sun et al., 2023). However, extending MBRL to multi-robot cooperation remains challenging. Early decentralized model-based multi-agent reinforcement learning (MARL) approaches built independent world models for each agent (Egorov & Shpilman, 2022), overlooking inter-agent couplings and hindering coordination. More recent centralized methods, by contrast, assume full observability or unrestricted communication (Chai et al., 2024; Toledo, 2024; Liu et al., 2024b), performing dynamics modeling and policy optimization in the joint space. These methods face challenges related to modeling complexity and synchronous communication in robotic systems with high-dimensional observation and action spaces, limiting their deployment in real-world scenarios.
Between centralized and decentralized paradigms, the distributed sequential paradigm has rapidly developed in recent years and demonstrated unique advantages (Khan, 2025). It reformulates multi-agent decision-making as an autoregressive process: agents communicate and act in a certain order, with each updating its policy conditioned on messages and actions from predecessors (Wen et al., 2022; Hu et al., 2025). This design enables more consistent joint reasoning (Ding et al., 2024) and finer-grained credit assignment (Kuba et al., 2022; Wang et al., 2023) without relying on full communication. From a real-world deployment perspective, it reduces the reliance on communication synchronization and offers improved robustness against packet loss or disturbances (Ding et al., 2024).
Motivated by these advantages, as shown in Figure 1, we propose the Sequential World Model (SeqWM), which integrates the sequential paradigm into model-based MARL to structurally decompose dynamics modeling and action planning. For trajectory prediction, SeqWM represents the joint dynamics as sequential agent-wise rollouts following the communication order, where each agent maintains an independent world model and conditions on the predicted trajectories and actions of its predecessors. For action planning, each agent performs multi-step lookahead conditioned on its predecessors’ predictions, thereby preserving cooperative performance while constraining the search to a low-dimensional subspace consistent with the sequential structure. We evaluate SeqWM in two challenging multi-robot cooperation environments: Bi-DexHands (Chen et al., 2024a) and Multi-Quad (Xiong et al., 2024), and further validate its effectiveness on physical multi-robot tasks using two Unitree Go2-W robots. The key contributions are as follows:
@@
    By integrating the sequential paradigm, SeqWM decomposes joint dynamics into autoregressive agent-wise models, reducing modeling complexity and the need for synchronized communication, thereby extending model-based MARL to multi-robot cooperation.
@@
    Through explicit intention sharing, SeqWM enables the emergence of advanced cooperative behaviors such as predictive adaptation, temporal alignment, and role division.
@@
    SeqWM achieves state-of-the-art performance in both simulation and real-world quadruped experiments, notably becoming the first multi-agent method to succeed on BottleCap and PushBox.
These results collectively demonstrate that SeqWM, by leveraging sequential world modeling and planning, offers an effective pathway for multi-robot cooperation, balancing performance, efficiency, and real-world applicability.
@@ 2 Related Work
Model-based RL. In single-agent robotics, model-based RL has shown remarkable success due to its high sample efficiency (Zhou et al., 2024; Lancaster et al., 2024), with approaches such as PlanCP (Sun et al., 2023) and GPC (Qi et al., 2025) leveraging learned dynamics models to predict trajectories and optimize actions for physical robots. In contrast, existing model-based MARL methods often rely on centralized paradigms (Zhao et al., 2025b), hindering their practical deployment in multi-robot systems. For example, MACD (Chai et al., 2024) and CoDreamer (Toledo, 2024) use transformers or GNNs to integrate full state-action across all agents. Recent efforts such as MARIE (Zhang et al., 2025a) explored decentralized dynamics modeling, but still require communication at each prediction step for agent-wise aggregation. Different from these works, SeqWM assigns each agent an independent world model and predicts trajectories sequentially, which not only lowers modeling complexity but also substantially reduces the reliance on synchronous communication, thereby making it applicable to real-world scenarios.
Sequential Paradigm. Recent studies in MARL have highlighted the advantages of the sequential paradigm (Khan, 2025), which enables fine-grained credit assignment (Kuba et al., 2022), efficient dynamics modeling (Zhang et al., 2025b), and scalable coordination (Xu et al., 2025). For example, MAT (Wen et al., 2022) and PMAT (Hu et al., 2025) model the multi-agent decision-making process as a sequence prediction problem, employing transformers to autoregressively predict each agent’s actions. A2PO (Wang et al., 2023) and HARL (Liu et al., 2024a; Zhong et al., 2024) further introduce the sequential update scheme that bring clearer interpretability and ensure monotonic improvement. SeqComm (Ding et al., 2024) extends this idea to the communication domain, where agents exchange information in a sequential order, effectively mitigating non-stationarity. Motivated by these benefits, we integrate the sequential paradigm into model-based MARL, leveraging world models to enhance planning and coordination.
@@ 3 Preliminaries
Problem Formulation. We model the fully cooperative task as a decentralized partially observable Markov decision process (dec-POMDP) (Zhao et al., 2025a), $\mathcal{M}=\langle\mathcal{I},\mathcal{S},\mathcal{O},\mathcal{A},\Omega,\mathcal{P},\mathcal{R},\gamma\rangle$, where $\mathcal{I}=\{v^{1},\dots,v^{n}\}$ is the set of agents, $\mathcal{S}$ is the global state space, $\mathcal{O}=\prod_{i=1}^{n}O^{i}$ is the joint observation space, and $\mathcal{A}=\prod_{i=1}^{n}A^{i}$ is the joint action space. The observation function $\Omega:\mathcal{S}\times\mathcal{I}\rightarrow\mathcal{O}$ defines each agent’s perception of the environment, while the transition function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ specifies the environment dynamics. The reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ provides a shared scalar signal, and $\gamma$ is the discount factor. Each agent $v^{i}$ learns a local policy $\pi^{i}:O^{i}\rightarrow A^{i}$, which maps its observation $o^{i}$ to an action $a^{i}$. The objective is to learn a joint policy $\pi=\prod_{i=1}^{n}\pi^{i}$ that maximizes the expected discounted return $\sum_{\tau=t}^{\infty}\gamma^{\tau}r_{\tau}$.
@@ 4 Methodology
In this section, we propose SeqWM, a Sequential World Model for multi-robot cooperation. Unlike existing centralized or decentralized multi-agent world models, SeqWM decomposes the joint dynamics into agent-wise models arranged in a sequence. This design substantially reduces modeling complexity and the reliance on synchronous communication, enabling deployment in physical multi-robot systems.
@@ 4.1 Sequential Multi-Agent World Model
Multi-Agent World Model. As noted in Introduction, multi-robot cooperative tasks involve high-dimensional observation and action spaces (Zhu et al., 2025a), making it unsuitable to use reconstructing raw observations as the learning objective of the world model (Hafner et al., 2025). Therefore we remove the explicit decoder and instead perform dynamics prediction entirely in a latent space. To facilitate distributed deployment, each agent maintains an independent world model without parameter sharing. Let $z^{i}_{t}$ denote the latent state of agent $v^{i}$ at timestep $t$, the world model can be defined as follows:
@@
All modules in SeqWM are implemented using MLPs, ensuring architectural simplicity and consistency. For communication function, we adopt concatenation operator $\oplus$ to facilitate modular training.111 We implement the concatenation using a mask-based scheme to ensure consistent dimensionality, with details provided in Appendix A.1. As shown in Section 5.5, this concise design achieves more stable training performance compared to alternatives such as cross-attention and recurrent neural networks (RNNs).
@@
where $\theta=\{\theta_{E},\theta_{D},\theta_{R},\theta_{Q}\}$, $H$ is the prediction horizon, $\lambda\in(0,1]$ is a constant that balances the contribution of each rollout step, $r_{t}$ is the ground-truth reward, $G_{t}$ is TD target, and $\hat{z}^{i}_{t+1}=D^{i}\left(z^{i}_{t},a^{i}_{t},e^{i}_{t}\right)$ is the predicted latent state. The loss can be backpropagated to the encoder via $z^{i}_{t}=E^{i}\left(o^{i}_{t}\right)$, so do the dynamics and reward losses. The latent target $z^{i}_{t+1}=E^{i}\left(o^{i}_{t+1}\right)$ is detached with the stop-gradient operator $\textrm{sg}(\cdot)$ to prevent cyclic gradient flow. Soft Cross-Entropy loss is used to match the discretized reward and Q-value predictions, with details provided in Appendix B.3. Additionally, to ensure modularity and scalability, each agent’s world model is trained independently, and the loss can not be backpropagated through the communication channel.
Based on Eq. (3), the encoder learns a compact latent space, while the dynamics and reward predictors minimize prediction errors in this space, ensuring alignment with real environment dynamics. The actor generates initial action estimates in the latent space to warm-start planning and is trained using the Heterogeneous-Agent Soft Actor-Critic (HASAC) (Liu et al., 2024a) algorithm:
@@
where $\alpha$ is the entropy coefficient, and $\mathcal{H}[\cdot]$ denotes the entropy function.
@@ 4.2 Planning with Sequential Communication
Although Eq. (2) includes an actor, it does not serve directly as an explicit decision policy; instead, it provides initial action estimates for the planner. We next propose a sequential multi-agent planner based on Model Predictive Path Integral (MPPI) (Williams et al., 2015) that leverages the predictions of world models to optimize each agent’s action. In this framework, the actor contributes by narrowing the action search space to promising regions, while the planner ensures robust long-term decision-making and corrects suboptimal proposals from the actor.
$V^{i}_{t+H}=\gamma^{H}Q^{i}(\hat{z}^{i}_{t+H},a^{i}_{t+H},e^{i}_{t+H})+\sum_{h=t}^{t+H-1}\gamma^{h-t}R^{i}(\hat{z}^{i}_{h},a^{i}_{h},e^{i}_{h}),$ (5)
@@
where $V^{i}_{t+H}$ represents the value estimate of a sampled action sequence, computed as the sum of predicted rewards over the horizon plus the terminal value given by the critic. Then, candidate sequences are ranked according to their evaluated values, and the highest-scoring subset is selected as the elite set. The action distribution is updated toward the statistics of these elite trajectories, thereby concentrating future sampling around high-value regions and progressively refining the search space. 222 The details of the planner and its implementationare provided in Appendix A.1 and  B.
@@
Low-pass Action Smoothing. To prevent mechanical wear caused by abrupt action changes, we integrate a low-pass filtering strategy (Kicki, 2025). In each planning iteration, sampled action sequences are filtered along the temporal dimension to suppress high-frequency fluctuations. This smoothing enforces gradual action transitions across timesteps, reducing control discontinuities and promoting stable, consistent behavior on physical robots. Further details are provided in Appendix A.2.
Heuristic Early-Stopping. Considering the computational constraints of physical robotic platforms, we design a motion-planning heuristic that terminates iterations when the KL (Kullback & Leibler, 1951) divergence between consecutive action distributions falls below a threshold. This early-stopping criterion mitigates diminishing returns (Kobilarov, 2012), reducing computation while preserving plan quality. Further details and experiments regarding this design are provided in Appendix C.5.